cdp:

Bill performed an experiment in which he spun a spinner divided into four equal parts. Each part had a different color. The results are shown below. Color Times Spun red 10 blue 17 green 20 yellow 13 Based on the experiment above, predict the number of times the arrow will stop on the yellow part of the spinner if he performs the experiment a total of 180 times. A. 63 B. 117 C. 126 D. 39

NotLeiland:

10+17+20+13 = 60. 180/60=3 3 x 13 =???

NotLeiland:

@notleiland wrote:
10+17+20+13 = 60. 180/60=3 3 x 13 =???
Your answer is 3 times 13